write the equation of the line parallel to the line give by 3x-2y=8 and having the same y intercept as the line given by 2x+3y=-4
Found 2 solutions by lwsshak3, nerdybill:
Answer by lwsshak3(11628)

write the equation of the line parallel to the line give by 3x-2y=8 and having the same y intercept as the line given by 2x+3y=-4
***
3x-2y=8
2y=3x-8
y=3x/2-4
slope=3/2
..
2x+3y=-4
3y=-2x-4
y=-2x/3-4/3
..
slope of line=3/2
y-intercept=-4/3
equation: y=3x/2-4/3

Answer by nerdybill(7384)

write the equation of the line parallel to the line give by 3x-2y=8 and having the same y intercept as the line given by 2x+3y=-4
.
slope of:
3x-2y=8
3x = 2y+8
3x-8 = 2y
(3/2)x - 4 = y
slope (because of "slope-intercept" form) is:
m = 3/2
The new line, since it is parallel, must also have a slope of 3/2
.
y-intercept of:
2x+3y=-4
3y = -2x-4
y = (-2/3)x - 4/3
y-intercept is -4/3
.
new line has slope of 3/2 and y-intercept of -4/3
plug into "slope-intercept" form:
y = mx + b
y = (3/2)x - 4/3 (answer)